# TDHook: A Lightweight Framework for Interpretability

### v1

@@ 5.2 Complex Models
While the previous example showcased a shared backbone for the policy and the value, it is not always the case. Especially in decentralised multi-agent settings, the value is often independent from the policy, an each agent can have its own policy and value. Therefore, it means that you have multiple networks to explain. Yet, in TorchRL, these models are grouped in a single TensorDictModule, the loss, which can be directly wrapped in TDHook. For our illustration, we train an agent with PPO (Schulman et al., 2017) on the inverted double pendulum environment (Todorov et al., 2012). More details about the agent studied are given in Appendix B. In this case, we decided to explore how the predicted action is represented throughout the internal states of this agent using linear probing (Alain & Bengio, 2018). Figure 10 shows the action probing in the agent, performed both on the policy and the value by wrapping the loss module. For each layer (pre- and post-activation), a linear regression probe is trained to predict the model’s action. We can see that the policy head contains more information about the action than the value head, as expected, but the value head still overperforms the baseline trained from the raw observations. For the policy, the last layers contain less information about the predicted action because of the compression and randomness of actions, since the action is sampled from a distribution. For the value, the information about the action is not trackable in the last layers, and the probe overfitted on the training data.
@@ 6.3 Future Work
Future releases of TDHook will focus on three complementary directions. First, we will broaden the set of integrated methods by porting additional attribution, probing and causal-manipulation techniques, with an emphasis on more complex pipelines. Second, we plan to exploit the advanced features of tensordict, in particular memory-mapped tensors and zero-copy sharing, to further reduce the peak RAM footprint during large-scale analyses. Lastly, we aim to extend support to distributed and heterogeneous execution environments so that TDHook can seamlessly operate in multi-GPU and multi-node training loops. These technical improvements will be accompanied by richer documentation and end-to-end tutorials that showcase the library on vision, language and control tasks, with primary focus on composed interpretability pipelines.
@@
As an interpretability tool, the framework proposed does not raise immediate ethical risks beyond those inherent to interpretability research. However, the goal being to make interpretability methods easier to use and more accessible to non-experts, it can exacerbate misuse cases. The most concerning issues are:
@@ Appendix B Use Cases Details
For our analysis, we choose the VGG16 model (Simonyan & Zisserman, 2014), loaded using the timm library (Wightman, ). We train linear probes for concepts on the texture dataset (Cimpoi et al., 2013), using the scikit-learn library (Pedregosa et al., 2011). As pointed out in (Dreyer et al., 2023b), the difference of means (or signal CAV (Pahde et al., 2022)) is more effective to find latent concept directions, as shown in figures 16 and 17.
@@
Appendix C Codebase Outline
@@
Attribution 13 Simonyan et al. (2013) Shrikumar et al. (2016) Sundararajan et al. (2017) Dhamdhere et al. (2018) Selvaraju et al. (2016) Bach et al. (2015) Lapuschkin et al. (2019) Montavon et al. (2019) Andéol et al. (2021) Montavon et al. (2015) Achtibat et al. (2022) Mahendran & Vedaldi (2015) Springenberg et al. (2014)
Latent Manipulation 8 Chen et al. (2020) Abnar & Zuidema (2020) Alain & Bengio (2018) Kim et al. (2018) Vig et al. (2020) Belrose et al. (2023) Dreyer et al. (2023b) Rimsky et al. (2023)
Weights 6 Meng et al. (2022) Cunningham et al. (2023) Dunefsky et al. (2024) Ilharco et al. (2022) Yeom et al. (2019) Pochinkov & Schoots (2024)
@@
  • activation_maximisation.py: activation maximisation (Mahendran & Vedaldi, 2015).
@@
  • activation_caching.py: maximally activating samples (Chen et al., 2020) and attention visualisation (Abnar & Zuidema, 2020).
  • probing.py: linear probing (Alain & Bengio, 2018) and concept activation vectors (Kim et al., 2018).
@@
  • pruning.py: relevance-based pruning (Yeom et al., 2019) and circuit pruning (Pochinkov & Schoots, 2024).
Future work could extend attribution with occlusion-based methods (Zeiler & Fergus, 2013), latent activation with AtP* (Kram’ar et al., 2024), or weights-based methods with crosscoders that generalise adapters to multiple layers and models (Lindsey et al., 2024).

### v2

@@ 5.2 Complex Models
While the previous example showcased a shared backbone for the policy and the value, it is not always the case. Especially in decentralised multi-agent settings, the value is often independent from the policy, and each agent can have its own policy and value. Therefore, it means that you have multiple networks to explain. Yet, in TorchRL, these models are grouped in a single TensorDictModule, the loss, which can be directly wrapped in TDHook. For our illustration, we train an agent with PPO (Schulman et al., 2017) on the inverted double pendulum environment (Todorov et al., 2012). More details about the agent studied are given in Appendix B. In this case, we decided to explore how the predicted action is represented throughout the internal states of this agent using linear probing (Alain and Bengio, 2018). Figure 10 shows the action probing in the agent, performed both on the policy and the value by wrapping the loss module. For each layer (pre- and post-activation), a linear regression probe is trained to predict the model’s action. We can see that the policy head contains more information about the action than the value head, as expected, but the value head still overperforms the baseline trained from the raw observations. For the policy, the last layers contain less information about the predicted action because it is highly compressed, only predicting the mean and standard deviation of a Tanh-Normal distribution. For the value, the information about the action is not trackable in the last layers, and the probe overfitted on the training data, as it can be expected when the signal weakens.
@@ 6.3 Future Work
Future releases of TDHook will focus on three complementary directions. First, we will broaden the set of integrated methods by porting additional attribution, probing and causal-manipulation techniques, with an emphasis on more complex pipelines. This includes extending our work to more sophisticated models, like transformers with chain-of-thought modules, requiring large-scale analyses. Second, we plan to exploit the advanced features of tensordict, in particular memory-mapped tensors and zero-copy sharing, to further reduce the peak RAM footprint during large-scale analyses. Lastly, we aim to extend support to distributed and heterogeneous execution environments so that TDHook can seamlessly operate in multi-GPU and multi-node training loops. These technical improvements will be accompanied by richer documentation and end-to-end tutorials that showcase the library on vision, language and control tasks, with primary focus on composed interpretability pipelines.
@@
As an interpretability tool, the framework proposed does not raise immediate ethical risks beyond those inherent to interpretability research. However, the goal of making interpretability methods easier to use and more accessible to non-experts can exacerbate misuse cases. The most concerning issues are:
@@ Appendix B Use Cases Details
For our analysis, we choose the VGG16 model (Simonyan and Zisserman, 2014), loaded using the timm library (Wightman, ). We train linear probes for concepts on the texture dataset from (Cimpoi et al., 2013), using the scikit-learn library (Pedregosa et al., 2011). As pointed out in (Dreyer et al., 2023b), the difference of means (or signal CAV (Pahde et al., 2022)) is more effective to find latent concept directions, as shown in figures 16 and 17.
@@
Appendix C Simple Use Cases
This appendix showcases simple, self-contained examples of using TDHook for common interpretability tasks.
Attribution
TDHook provides a unified API for attribution methods, making it simple to switch between different techniques. In Listing C.1, we demonstrate how to compute attribution maps for a pre-trained VGG16 model. We highlight the minimal changes required to switch from Saliency (red lines) to IntegratedGradients (green lines), and present both results in Figure 20.
1import torch
2import timm
3from PIL import Image
4from tensordict import TensorDict
5from tdhook.attribution import Saliency
6from tdhook.attribution import IntegratedGradients
7
8# Load model and prepare image
9model = timm.create_model("vgg16.tv_in1k", pretrained=True)
10data_config = timm.data.resolve_model_data_config(model)
11transforms = timm.data.create_transform(**data_config, is_training=False)
12
13image = Image.open("results/simple/zebra_1.jpg").convert("RGB")
14image_tensor = transforms(image)
15
16# Define attribution target (zebra class = 340)
17def init_attr_targets(targets, _):
18 zebra_logit = targets["output"][..., 340]
19 return TensorDict(out=zebra_logit, batch_size=targets.batch_size)
20
21# Compute attribution
22with Saliency(
23with IntegratedGradients(
24 init_attr_targets=init_attr_targets
25).prepare(model) as hooked_model:
26 td = TensorDict(
27 {
28 "input": image_tensor.unsqueeze(0),
29      ("baseline", "input"): torch.zeros_like(image_tensor).unsqueeze(0)
30 },
31 batch_size=1,
32 )
33 td = hooked_model(td) # Access attribution with td.get(("attr", "input"))
Listing C.1: Code example showing the switch from Saliency to IntegratedGradients.
Steering Vectors
TDHook supports intervention techniques such as steering vectors which have proven effective for language models (Rimsky et al., 2023). In Listing C.2, we demonstrate how to extract a steering vector from GPT-2 that represents the concept of "wealth" and use it to steer the model’s generation. When completing the prompt "I work as a", the model is steered from "writer" to "pilot", reflecting the injected concept [^4] Additional samples could be used to extract more robust and representative steering vectors..
1from transformers import AutoTokenizer, AutoModelForCausalLM
2from tensordict import TensorDict
3from tdhook.latent import ActivationAddition, SteeringVectors
4
5# Load model and tokenizer
6model = AutoModelForCausalLM.from_pretrained("gpt2")
7tokenizer = AutoTokenizer.from_pretrained("gpt2")
8
9# Prepare inputs
10positive_inputs = tokenizer.encode("I am rich.", return_tensors="pt")
11negative_inputs = tokenizer.encode("I am poor.", return_tensors="pt")
12base_inputs = tokenizer.encode("I work as a", return_tensors="pt")
13
14# Extract steering vector (rich - poor)
15with ActivationAddition(["transformer.h.7.mlp"]).prepare(model) as hooked_model:
16 td = TensorDict(
17 {
18 ("positive", "input"): positive_inputs,
19 ("negative", "input"): negative_inputs
20 },
21 batch_size=1
22 )
23 td = hooked_model(td)
24
25steering_vector = td.get(("steer", "transformer.h.7.mlp")).sum(dim=0)
26
27
28# Define steering function
29def steer_fn(module_key, output):
30 return output + 4 * steering_vector
31
32
33# Apply steering during inference
34with SteeringVectors(["transformer.h.7.mlp"], steer_fn=steer_fn).prepare(model) as hooked_model:
35 td = TensorDict({"input": base_inputs}, batch_size=1)
36 td = hooked_model(td)
37
38# Compare results
39steered_token = td.get(("output", "logits")).max(dim=-1).indices[0, -1]
40original_token = model(base_inputs)["logits"].max(dim=-1).indices[0, -1]
41
42print(f"Steered: {tokenizer.decode(steered_token)}") # Output: "pilot"
43print(f"Original: {tokenizer.decode(original_token)}") # Output: "writer"
Listing C.2: Code example showing the extraction and application of a steering vector in GPT-2.
(a) Original image
(b) Saliency
(c) Integrated Gradients
Figure 20: Comparison of attribution maps for the "zebra" class (index 340) using VGG16. Saliency (left) and Integrated Gradients (right) are computed using the code above.
Appendix D Codebase Outline
@@
Attribution 13 Simonyan et al. (2013) Shrikumar et al. (2016) Sundararajan et al. (2017) Dhamdhere et al. (2018) Selvaraju et al. (2016) Bach et al. (2015) Lapuschkin et al. (2019) Montavon et al. (2019) Andéol et al. (2021) Montavon et al. (2015) Achtibat et al. (2022) Mahendran and Vedaldi (2015) Springenberg et al. (2014)
Latent Manipulation 8 Chen et al. (2020) Abnar and Zuidema (2020) Alain and Bengio (2018) Kim et al. (2018) Vig et al. (2020) Belrose et al. (2023) Dreyer et al. (2023b) Rimsky et al. (2023)
Weights 6 Meng et al. (2022) Cunningham et al. (2023) Dunefsky et al. (2024) Ilharco et al. (2022) Yeom et al. (2019) Pochinkov and Schoots (2024)
@@
  • activation_maximisation.py: activation maximisation (Mahendran and Vedaldi, 2015).
@@
  • activation_caching.py: maximally activating samples (Chen et al., 2020) and attention visualisation (Abnar and Zuidema, 2020).
  • probing.py: linear probing (Alain and Bengio, 2018) and concept activation vectors (Kim et al., 2018).
@@
  • pruning.py: relevance-based pruning (Yeom et al., 2019) and circuit pruning (Pochinkov and Schoots, 2024).
Future work could extend attribution with occlusion-based methods (Zeiler and Fergus, 2013), latent activation with AtP* (Kram’ar et al., 2024), or weights-based methods with crosscoders that generalise adapters to multiple layers and models (Lindsey et al., 2024).